So you want to be an astronaut. There is a new “test” that all must go through call the vertical circle of death (well at least that is what we will call it). On this “ride” you (m=76kg) are spun in a 18m diameter vertical circle once every 1.5s.

a) what is your angular speed?

b) what is the magnitude of your centripetal acceleration?

c) what is your apparent weight at the top and the bottom of the ride?
Answer #1

a)

w = 2 pi / T = 2* 3.14 / 1.5 = 4.187 rad/s

b)

ac = r w^2 = 9* 4.187^2 = 157.75 m/s^2

c)

apparent weight at top

W = m ac - mg

W = 76* ( 157.75 - 9.8)

W = 11244.5 N

apparent weight at bottom

W = 76 ( ac + g)

W = 12733.8 N

==========

Comment in case any doubt, will reply for sure.. Goodluck